# Modulate and Reconstruct: Learning Hyperspectral Imaging from Misaligned Smartphone Views

### v3

@@ 8 Conclusion
This work rethinks HSR through the lens of practical acquisition. By moving beyond single RGB image constraints, we demonstrate that leveraging multiple smartphone cameras with carefully chosen spectral filters can significantly enrich the input signal and reduce ambiguity in reconstruction. Our proposed MI-HSR framework shows that spectral diversity even when captured with commodity hardware can close the gap between simulated and real-world hyperspectral imaging. The introduction of the Doomer dataset marks an important step forward for benchmarking in this space, enabling systematic evaluation of multi-view HSR under realistic conditions. Empirical results validate both our hardware configuration and model design, suggesting a promising direction for low-cost, deployable hyperspectral imaging systems. Looking ahead, we aim to further explore the temporal dimension for dynamic scenes, optimize for energy-efficient mobile deployment, and investigate more principled learning paradigms under limited supervision or device mismatch.
@@ Appendix A Technical Appendices and Supplementary Material
  • $\mathbf{c}$ — a photometrically normalized signal (probably noisy) from smartphone camera(s). Can be $\in\mathbb{R}^{3}$ or $\in\mathbb{R}^{3k}$ depending on context.
  • $\bar{\mathbf{c}}$ — a photometrically normalized signal as it would be in an ideal world without noise.
  • $\mathbf{r}$ — a photometrically normalized radiance spectrum. Is assumed to be noiseless (however, in Sec. 6.4 of the main paper we discuss the outcomes of this assumption).
  • $\mathbf{x}_{i}$ — a vector from a set of vectors $\{\mathbf{x}_{1},\mathbf{x}_{2},$...\}
  • $x_{i}$$i$-th component of $\mathbf{x}$.
@@ A.1 Estimation of Spectral Sensitivity Functions and Transmittance Functions
Spectral sensitivity functions. To estimate SSF of a smartphone camera, we acquired 25 sample pairs of $\mathbf{c}_{i}\in\mathbb{R}^{3}$ and $\mathbf{r}_{i}\in\mathbb{R}^{n}$ corresponding to flat-field illumination (FFI) of narrow-band LED light sources in an integrating sphere. For each LED $i$, we took a photo of FFI (Fig. 8), extracted the central 100×100 patch from it, photometrically normalized it and computed the channel-wise average to get $\mathbf{c}_{i}$. Then we measured the LED radiance spectrum $\mathbf{r}_{i}$ using an X-Rite i1 Pro spectrophotometer with the help of spotread routine from Argyll color management system.
@@
Figure 8: Training data for SSF estimation: (a) spectra of 25 LEDs measured using X-Rite i1 Pro; (b) an example of raw FFI photo. Notice the vignetting effect. To mitigate this effect, channel-wise average of a tiny central patch (in red) was used as a sample $\mathbf{c}_{i}$
Given the measurements $\{\mathbf{c}_{i},\mathbf{r}_{i}\}_{i=1}^{25}$, the spectral sensitivities estimation problem can be formulated as a regularized quadratic optimization problem:
$\begin{aligned} \min_{\mathbf{S}}&\quad\sum_{i=1}^{25}\|\mathbf{c}_{i}-\mathbf{S}^{\top}\mathbf{r}_{i}\|_{2}^{2}+\lambda\|\mathbf{DS}\|_{2}^{2}\\ \text{s. t.}&\quad\mathbf{S}\geq 0\end{aligned}$ (1)
where the regularization term $\lambda\|\mathbf{DS}\|_{2}^{2}$ imposes smoothness, $\mathbf{D}$ is the second-order derivative operator. The objective (1) was minimized using the Adam optimizer.
@@
To validate the estimated $\hat{\mathbf{S}}$, we captured a color rendition chart by the Specim IQ hyperspectral camera and all three smartphone cameras. The HSI was calibrated (more on that in the next paragraph) and projected onto each camera’s sensor space to yield RGB predictions. We compared the predicted and actual RGBs of the color patches: pixels of a patch were pixel-wise averaged (Fig. 12 gives a sight on how such patches look like). The average angular error of color reproduction was $\approx 1^{\circ}$ — indistinguishable to a human eye.
Specim IQ calibration. Spectral measurements of X-Rite i1 Pro and Specim IQ are actually inconsistent. It is not suprising: Specim IQ is not designed to be photometrically accurate. Instead, it allows to estimate band-wise relations between the calibration target in a scene and other objects. However, if we estimate the band-wise calibration divisor between X-Rite i1 Pro and Specim IQ, we can futher use it to obtain physically-correct HSIs. We illuminated the integration sphere using a mixture of LEDs that yields even spectral power distribution, obtained X-Rite and Specim IQ measurements of it and divided one by another:
$\mathbf{k}=\mathbf{r}_{\mathrm{specim}}\oslash\mathbf{r}_{\mathrm{xrite}},$
@@ A.2 Spectral Uncertainty
where $H\left(\xi\mid\eta=x\right)$ is the ordinary entropy of $\xi$ given event $\eta=x$. However, entropy is designed especially for discrete-distributed variables of categorical type. When uncertainty of a categorical random variable is indeed best described by the entropy of its distribution, for continuous random variables variance is a better choice. When dealing with random vectors, variance becomes a matrix, so we consider the trace of it as a reasonable summary. If we put the trace of variance instead of the entropy in (2) and substitute $\xi=\mathbf{r},\eta=\mathbf{c}$, we get the spectral uncertainty of our filters-modified optical system:
$v(\mathbf{F})=\mathbb{E}_{\mathbf{c}}\left[\mathrm{tr}\,\mathbb{V}_{\mathbf{r}}(\mathbf{r}\mid\mathbf{c})\right],$
where the dependence on $\mathbf{F}$ is hidden inside the relationship between $\mathbf{r}$ and $\mathbf{c}$. However, $\mathbf{r}$ and $\mathbf{c}$ are not random variables until we define them so. Let $\mathbf{r}$ have a discrete distribution $p\left(\mathbf{r}_{i}\right)$ over a finite set of radiance spectra $\mathcal{R}=\{\mathbf{r}_{1},...,\mathbf{r}_{N}\}$ derived from a dataset of HSIs. Such definition reflects the a priori information about natural radiance. In the relation $\mathbf{c}=\mathbf{S_{F}r+n}$ only noise $\mathbf{n}$ is yet to define. Let $\bar{\mathbf{c}}=\mathbf{S_{F}r}$ be an unnoised camera response. We model noise as
$\mathbf{n}\mid\bar{\mathbf{c}}\;\sim\;\mathcal{N}(\mathbf{0},\mathbf{\Sigma(\bar{\mathbf{c}})}),$ (3)
$\mathbf{\Sigma}(\bar{\mathbf{c}})=\mathrm{diag}[\sigma^{2}_{1}(\bar{c}_{1});...;\sigma^{2}_{3k}(\bar{c}_{3k})],$ (4)
@@
Now, when we defined $p\left(\mathbf{c}\mid\mathbf{r}\right)$, we can derive from the Bayesian rule:
$p\left(\mathbf{r}_{i}\mid\mathbf{c}\right)=\frac{p\left(\mathbf{c}\mid\mathbf{r}_{i}\right)p\left(\mathbf{r}_{i}\right)}{\sum_{i=1}^{N}p\left(\mathbf{c}\mid\mathbf{r}_{i}\right)p\left(\mathbf{r}_{i}\right),}$
@@
$\mathbb{E}_{\mathbf{r}}\left(\mathbf{r}\mid\mathbf{c}\right)=\sum_{i=1}^{N}p\left(\mathbf{r}_{i}\mid\mathbf{c}\right)\mathbf{r}_{i},$
$\mathrm{tr}\,\mathbb{V}_{\mathbf{r}}\left(\mathbf{r}\mid\mathbf{c}\right)=\sum_{i=1}^{N}p\left(\mathbf{r}_{i}\mid\mathbf{c}\right)\left\|\mathbf{r}_{i}-\mathbb{E}_{\mathbf{r}}\left(\mathbf{r}\mid\mathbf{c}\right)\right\|^{2}_{2}.$
@@
Now we have an expression depending on $\mathbf{c}$ which we should take $\mathbb{E}_{\mathbf{c}}$ of. However, this is intractable. So we should resort to Monte Carlo method by sampling $\mathbf{c}$ according to its definition: first take random $\mathbf{r}^{*}\in\mathcal{R}$, then sample $\mathbf{c}$ from $\mathcal{N}(\mathbf{S_{F}r^{*}},\Sigma(\mathbf{S_{F}r^{*}}))$. In our experiments, $2^{20}$ samples were sufficient to achieve 0.5% of relative standard deviation.
@@
To derive $\mathcal{R}$, we first extracted every 29×29th pixel of the KAUST dataset [39], which resulted in 559,921 samples. However, this is a dataset of reflectances, so we converted them to radiance spectra via multiplying it by a spectrum of gray ball under overcast weather condition. This way, we assured that $\mathcal{R}$ mostly fairly represents our future Doomer dataset. To reduce computational requirements, we also compressed $\mathcal{R}$ to the size of $N=1024$ by running K-Means algorithm and employing clusters’ centers. We assigned $p\left(\mathbf{r}_{i}\right)$ to be the share of initial 559,921 radiance spectra that belong to the cluster around $\mathbf{r}_{i}$.
Unfortunately, our initial implementation of spectral uncertainty contained a bug in the code, so the selected filters pair (Fig. 2) is not actually the best according to the described criterion. The bug has been spot long after the dataset was collected. In Fig. 11 we show the best filters selected by the criterion after the bug fix. We also probe those filters in simulated settings in Sec. A.5
Figure 11: Best filters after the bug fix
@@
Noise model. Let $\bar{c}_{i}$ be a photometrically-normalized ground-truth signal level at channel $i$, $t$ be exposure time, $g$ be ISO. The expected collected charge at image sensor is $\bar{c}_{i}t$. Following Foi et al. [44], we model noise of the collected charge as $\varepsilon\sim\mathcal{N}(0,\alpha_{i}\bar{c}_{i}t+\beta_{i})$, where $\alpha_{i}$, $\beta_{i}$ are parameters of the model. The final value in the raw image is further amplified by $g$ along with the noise: $(\bar{c}_{i}t+\varepsilon)g$. Foi et al. also introduce additional noise after amplification, but we neglect it for simplicity. The observed photometrically-normalized signal is then given by:
$c_{i}=\frac{(\bar{c}_{i}t+\varepsilon)g}{tg}=\bar{c}_{i}+\frac{\varepsilon}{t}.$
@@
$\sigma_{i}(\bar{c}_{i})=\frac{\sqrt{\alpha_{i}\bar{c}_{i}t+\beta_{i}}}{t}$ (5)
We estimate parameters $\alpha_{i}$, $\beta_{i}$ of this model by plotting $\sigma_{i}t$ against $\bar{c}_{i}t$ in Fig. 12.
Figure 12: Plotting $\sigma_{i}t$ against $\bar{c}_{i}t$ for the Telephoto camera based on a single shot. Each data point (24 in total) is given by channel-wise mean and std dev of a color patch (left, in red). In each plot, point cloud is approximated by a dashed line (5) using least-squares; each color channel $i$ gets its own noise parameters $\alpha_{i}$, $\beta_{i}$ (right).
The only problem left in (5) is the hidden dependence on $t$. Given simulated $\bar{\mathbf{c}}=\mathbf{S_{F}r}$, we cannot add noise properly unless we know exposure time $t$. In real-life setting, exposure time is determined by a camera automatically depending on the brightness of a scene. For simplicity, we model this in a piecewise-power fashion. We gathered 1000 raw images from phone’s gallery and plotted $t$ against the average of photometrically normalized signals across all channels, height and width in an image $c=\frac{1}{h\times w\times 3}\sum_{i,j,k}\mathbf{I}(i,j,k)$ (Fig. 13). The fitted function allows to estimate $t$ given $\bar{c}_{i}$.
@@ A.3 More on Doomer Dataset
Both the single- and multi-camera versions will also have two subversions depending on what spectral resolution of an HSI a user needs. Our resources allowed us to capture spectral radiance with the sampling interval of 3 nm. However, most of hyperspectral images have that of 10 nm as a compromise between precision and computational efficiency. A short summary of each version provided in Tab. 6. At the time of submission, only emphasized items are made available.
The preprocessing pipeline of RAW images consists of these steps: demosaicing, black current subtraction, flat field calibration and photometric normalization. For hyperspectral images, preprocessing pipeline includes: black current subtraction, flat field calibration, radial distortion correction and photometric normalization. The last step involves only division by the exposure time (Specim IQ has no ISO setting) and band-wise division by the calibration divisor $\mathbf{k}$.
@@
A.4 Rest of the MI-MSFN Architecture
Since the proposed neural network is rather a minor contribution and differs from MSFN only by the arrangement of blocks, we only gave a broad view on it in the main part of the paper. Some blocks were left unexpanded, which we fix by showing the missing details here in Fig. 14.
Figure 14: Illustration of fusion modules and feed-forward network in MI-MSFN. GAP and GMP stand for global average and max pooling respectively.
The remaining disambiguation consists of abbreviations expansion: S-MSA = spectral multihead self-attention [8]; SW-MSA = shifted window multihead self-attention [46]; W-MSA = window multihead self-attention [47].
A.5 Additional Experiments on Simulated Data
Having the estimated spectral characteristics of our filtered smartphone cameras setup ($\mathbf{S_{F}}$) and the sensor noise model (5), we find ourselves in a position, where we can simulate images of the triple-camera setup from HSIs and run experiments on simulated data. Unlike the real-life conditions, the simulated images are perfectly aligned to HSI and between each other. Evaluation on a simulated dataset then gives us a sight on how big the performance increase caused by the triple-camera setup could be if there was no misalignment at all. The results are shown in Tab. 7
Setting PSNR, dB ↑ SAM, ° ↓ NSE, % ↓
Single-camera 28.79 ± 0.46 5.40 ± 0.22 8.46 ± 0.37
Multi-camera 36.84 ± 0.39 2.69 ± 0.02 3.88 ± 0.06
Multi-camera (filters from Fig. 11) 37.90 ± 0.31 2.33 ± 0.06 3.43 ± 0.08
Table 7: Evaluations of MI-MSFN in single- and multi-camera settings on simulated data. Evaluation of the better filters combination is provided separately
The triple-camera setup in simulated conditions yields $\sim 50\%$ improvement in terms of SAM and NSE and +8 dB PSNR. Recalling the $\sim 30\%$ improvement in the real-life conditions, we can conclude that our results in the main part of the paper are reasonable.
Tab. 7 also includes results for the filters selected after we fixed the bug in our code (Fig. 11). It suggests that if we used those filters for the dataset collection, we could have achieved even better results on the real data in the main paper. Also, this simulation result strengthens the reasonableness of spectral uncertainty criterion.
A.6 Computational Resources
All the network trainings (Tab. 2, 4, 4) were executed on a cluster, allowing us to run up to 20 jobs in parallel. This was especially handy because we needed to run each experiment 10 times with different random seeds to ensure statistic significance. The cluster randomly assigned jobs to nodes, each having multiple GPUs. However, any job only used one GPU at a time. A GPU assigned to a job could be one of the following: NVIDIA L40, L40s and H100. Every training consumed < 23 GB of GPU memory and lasted no longer than 8 hours. 5 CPU cores were occupied by data loader; the total CPU memory consumption was < 8 GB.

### v4

@@ 7 Conclusion
This work rethinks HSR through the lens of practical acquisition. By moving beyond single RGB image constraints, we demonstrate that leveraging multiple smartphone cameras with carefully chosen spectral filters can significantly enrich the input signal and reduce ambiguity in reconstruction.
The introduction of the Doomer dataset marks an important step forward for benchmarking in this space, enabling systematic evaluation of multi-view HSR under realistic conditions. Empirical results validate both our hardware configuration and model design, suggesting a promising direction for low-cost, deployable hyperspectral imaging systems. Proposed Deformable Convolution Alignment Module and conducted experiments show the alignment importance in the MI-HSR task, while proposed GMST++ is comparable with other state-of-the-art methods in established single image HSR it shows spectral reconstruction improvement in multi-image input even without DCAM.
Looking ahead, we aim to further explore the temporal dimension for dynamic scenes, optimize for energy-efficient mobile deployment, and investigate more principled learning paradigms under limited supervision or device mismatch.
@@ Appendix A Technical Appendices and Supplementary Material
  • $\mathbf{x}$ — a photometrically normalized signal (probably noisy) from smartphone camera(s). Can be $\in\mathbb{R}^{3}$ or $\in\mathbb{R}^{3k}$ depending on context.
  • $\bar{\mathbf{x}}$ — a photometrically normalized signal as it would be in an ideal world without noise.
  • $\mathbf{y}$ — a photometrically normalized radiance spectrum. Is assumed to be noiseless (however, in Sec. 6.4 of the main paper we discuss the outcomes of this assumption).
@@ A.1 Estimation of Spectral Sensitivity Functions and Transmittance Functions
Spectral sensitivity functions. To estimate SSF of a smartphone camera, we acquired 25 sample pairs of ${\bm{x}}_{i}\in\mathbb{R}^{3}$ and ${\bm{y}}_{i}\in\mathbb{R}^{n}$ corresponding to flat-field illumination (FFI) of narrow-band LED light sources in an integrating sphere. For each LED $i$, we took a photo of FFI (Fig. 6), extracted the central 100×100 patch from it, photometrically normalized it and computed the channel-wise average to get ${\bm{x}}_{i}$. Then we measured the LED radiance spectrum ${\bm{y}}_{i}$ using an X-Rite i1 Pro spectrophotometer with the help of spotread routine from Argyll color management system.
@@
Figure 6: Training data for SSF estimation: (a) spectra of 25 LEDs measured using X-Rite i1 Pro; (b) an example of raw FFI photo. Notice the vignetting effect. To mitigate this effect, channel-wise average of a tiny central patch (in red) was used as a sample ${\bm{x}}_{i}$
Given the measurements $\{{\bm{x}}_{i},{\bm{y}}_{i}\}_{i=1}^{25}$, the spectral sensitivities estimation problem can be formulated as a regularized quadratic optimization problem:
$\begin{aligned} \min_{\bm{C}}&\quad\sum_{i=1}^{25}\|{\bm{x}}_{i}-{\bm{C}}^{\top}{\bm{y}}_{i}\|_{2}^{2}+\lambda\|{\bm{D}}{\bm{C}}\|_{2}^{2}\\ \text{s. t.}&\quad{\bm{C}}\geq 0\end{aligned}$ (1)
where the regularization term $\lambda\|{\bm{D}}{\bm{C}}\|_{2}^{2}$ imposes smoothness, $\bm{D}$} is the second-order derivative operator. The objective (1) was minimized using the Adam optimizer.
@@
To validate the estimated $\hat{{\bm{C}}}$, we captured a color rendition chart by the Specim IQ hyperspectral camera and all three smartphone cameras. The HSI was calibrated (more on that in the next paragraph) and projected onto each camera’s sensor space to yield RGB predictions. We compared the predicted and actual RGBs of the color patches: pixels of a patch were pixel-wise averaged (Fig. 12 gives a sight on how such patches look like). The average angular error of color reproduction was $\approx 1^{\circ}$ — indistinguishable to a human eye.
Specim IQ calibration. Spectral measurements of X-Rite i1 Pro and Specim IQ are actually inconsistent. It is not suprising: Specim IQ is not designed to be photometrically accurate. Instead, it allows to estimate band-wise relations between the calibration target in a scene and other objects. However, if we estimate the band-wise calibration divisor between X-Rite i1 Pro and Specim IQ, we can further use it to obtain physically-correct HSIs. We illuminated the integration sphere using a mixture of LEDs that yields even spectral power distribution, obtained X-Rite and Specim IQ measurements of it and divided one by another:
${\bm{k}}={\bm{y}}_{\mathrm{specim}}\oslash{\bm{y}}_{\mathrm{xrite}},$
@@ A.2 Spectral Uncertainty
where $H\left(\xi\mid\eta=x\right)$ is the ordinary entropy of $\xi$ given event $\eta=x$. However, entropy is designed especially for discrete-distributed variables of categorical type. When uncertainty of a categorical random variable is indeed best described by the entropy of its distribution, for continuous random variables variance is a better choice. When dealing with random vectors, variance becomes a matrix, so we consider the trace of it as a reasonable summary. If we put the trace of variance instead of the entropy in (2) and substitute $\xi=\mathbf{y},\eta=\mathbf{x}$, we get the spectral uncertainty of our filters-modified optical system:
$v({\bm{F}})=\mathbb{E}_{\mathbf{x}}\left[\mathrm{tr}\,\mathrm{Var}_{\mathbf{y}}(\mathbf{y}\mid\mathbf{x})\right],$
where the dependence on $\bm{F}$} is hidden inside the relationship between $\mathbf{y}$ and $\mathbf{x}$. However, $\mathbf{y}$ and $\mathbf{x}$ are not random variables until we define them so. Let $\mathbf{y}$ have a discrete distribution $p\left(\mathbf{y}_{i}\right)$ over a finite set of radiance spectra $\mathcal{R}=\{\mathbf{y}_{1},...,\mathbf{y}_{N}\}$ derived from a dataset of HSIs. Such definition reflects the a priori information about natural radiance. In the relation $\mathbf{x}={\bm{C}}_{\bm{F}}\mathbf{r+n}$ only noise $\mathbf{n}$ is yet to define. Let $\bar{\mathbf{x}}={\bm{C}}_{\bm{F}}\mathbf{y}$ be an unnoised camera response. We model noise as
$\mathbf{n}\mid\bar{\mathbf{x}}\;\sim\;\mathcal{N}(\mathbf{0},\mathbf{\Sigma(\bar{\mathbf{x}})}),$ (3)
$\mathbf{\Sigma}(\bar{\mathbf{x}})=\mathrm{diag}[\sigma^{2}_{1}(\bar{c}_{1});...;\sigma^{2}_{3k}(\bar{c}_{3k})],$ (4)
@@
Now, when we defined $p\left(\mathbf{x}\mid\mathbf{y}\right)$, we can derive from the Bayesian rule:
$p\left(\mathbf{y}_{i}\mid\mathbf{x}\right)=\frac{p\left(\mathbf{x}\mid\mathbf{y}_{i}\right)p\left(\mathbf{y}_{i}\right)}{\sum_{i=1}^{N}p\left(\mathbf{x}\mid\mathbf{y}_{i}\right)p\left(\mathbf{y}_{i}\right),}$
@@
$\mathbb{E}_{\mathbf{y}}\left(\mathbf{y}\mid\mathbf{x}\right)=\sum_{i=1}^{N}p\left(\mathbf{y}_{i}\mid\mathbf{x}\right)\mathbf{y}_{i},$
$\mathrm{tr}\,\mathrm{Var}_{\mathbf{y}}\left(\mathbf{y}\mid\mathbf{x}\right)=\sum_{i=1}^{N}p\left(\mathbf{y}_{i}\mid\mathbf{x}\right)\left\|\mathbf{y}_{i}-\mathbb{E}_{\mathbf{y}}\left(\mathbf{y}\mid\mathbf{x}\right)\right\|^{2}_{2}.$
@@
Now we have an expression depending on $\mathbf{x}$ which we should take $\mathbb{E}_{\mathbf{x}}$ of. However, this is intractable. So we should resort to Monte Carlo method by sampling $\mathbf{x}$ according to its definition: first take random $\mathbf{y}^{*}\in\mathcal{R}$, then sample $\mathbf{x}$ from $\mathcal{N}({\bm{C}}_{\bm{F}}\mathbf{y}^{*},\Sigma({\bm{C}}_{\bm{F}}\mathbf{y}^{*}))$. In our experiments, $2^{20}$ samples were sufficient to achieve 0.5% of relative standard deviation.
@@
To derive $\mathcal{R}$, we first extracted every 29×29th pixel of the KAUST dataset [25], which resulted in 559,921 samples. To reduce computational requirements, we compressed $\mathcal{R}$ to the size of $N=1024$ by running K-Means algorithm and employing clusters’ centers. We assigned $p\left(\mathbf{y}_{i}\right)$ to be the share of initial 559,921 radiance spectra that belong to the cluster around $\mathbf{y}_{i}$.
The KAUST dataset contains objects reflectance. We correct work with radiance data we sample the illuminant spectra from gray ball after collecting the Doomer dataset. In Fig. 10 we show the best filters selected by the criterion after applying typical Doomer illumination.
Figure 9: Filters selected for Tele and Wide cameras.
Figure 10: Best filters for typical Doomer radiance.
Experimental justification. Having the estimated spectral characteristics of our filtered smartphone cameras setup (${\bm{C}}_{\bm{F}}$) and the sensor noise model (5), we find ourselves in a position where we can simulate images of the triple-camera setup from HSIs and run experiments on simulated data. Therefore, we can support the proposed spectral uncertainty criterion by simulating input images ${\bm{\mathsfit{X}}}_{0}$, ${\bm{\mathsfit{X}}}_{1}$, ${\bm{\mathsfit{X}}}_{2}$ given filters, measuring HSR performance and plotting it against $v({\bm{C}}_{\bm{F}})$.
We chose 12 random filter pairs and simulated 12 versions of our dataset. For each version, MI-MSFN was trained and evaluated. Fig. 11 shows the dependence of three performance metrics on the proposed spectral uncertainty criterion. A strong correlation is observed, as the Pearson correlation coefficient has an absolute value greater than 0.8.
(a) PSNR vs Variance
(b) SAM vs Variance
(c) NSE vs Variance
Figure 11: Relation between PSNR, SAM, NSE in simulated settings and the proposed spectral uncertainty criterion.
@@
Noise model. Let $\bar{x}_{i}$ be a photometrically-normalized ground-truth signal level at channel $i$, $t$ be exposure time, $g$ be ISO. The expected collected charge at image sensor is $\bar{x}_{i}t$. Following Foi et al. [13], we model noise of the collected charge as $\varepsilon\sim\mathcal{N}(0,\alpha_{i}\bar{x}_{i}t+\beta_{i})$, where $\alpha_{i}$, $\beta_{i}$ are parameters of the model. The final value in the raw image is further amplified by $g$ along with the noise: $(\bar{x}_{i}t+\varepsilon)g$. Foi et al. also introduce additional noise after amplification, but we neglect it for simplicity. The observed photometrically-normalized signal is then given by:
$x_{i}=\frac{(\bar{x}_{i}t+\varepsilon)g}{tg}=\bar{x}_{i}+\frac{\varepsilon}{t}.$
@@
$\sigma_{i}(\bar{x}_{i})=\frac{\sqrt{\alpha_{i}\bar{x}_{i}t+\beta_{i}}}{t}$ (5)
We estimate parameters $\alpha_{i}$, $\beta_{i}$ of this model by plotting $\sigma_{i}t$ against $\bar{x}_{i}t$ in Fig. 12.
Figure 12: Plotting $\sigma_{i}t$ against $\bar{x}_{i}t$ for the Telephoto camera based on a single shot. Each data point (24 in total) is given by channel-wise mean and std dev of a color patch (left, in red). In each plot, point cloud is approximated by a dashed line (5) using least-squares; each color channel $i$ gets its own noise parameters $\alpha_{i}$, $\beta_{i}$ (right).
The only problem left in (5) is the hidden dependence on $t$. Given simulated $\bar{\mathbf{x}}={\bm{C}}_{\bm{F}}\mathbf{y}$, we cannot add noise properly unless we know exposure time $t$. In real-life setting, exposure time is determined by a camera automatically depending on the brightness of a scene. For simplicity, we model this in a piecewise-power fashion. We gathered 1000 raw images from phone’s gallery and plotted $t$ against the average of photometrically normalized signals across all channels, height and width in an image $x=\frac{1}{h\times w\times 3}\sum_{i,j,k}{\bm{\mathsfit{X}}}_{i,j,k}$ (Fig. 13). The fitted function allows to estimate $t$ given $\bar{x}_{i}$.
@@ A.3 More on Doomer Dataset
Both the single- and multi-camera versions will also have two subversions of hyperspectral sampling interval: 3 nm or 10 nm. A short summary of each version provided in Tab. 6.
The preprocessing pipeline of RAW images consists of these steps: demosaicing, black current subtraction, flat field calibration and photometric normalization. For hyperspectral images, preprocessing pipeline includes: black current subtraction, flat field calibration, radial distortion correction and photometric normalization. The last step involves only division by the exposure time (Specim IQ has no ISO setting) and band-wise division by the calibration divisor $\bm{k}$}.
@@
The data preprocessing pipeline was described in Sec. 4 and is briefly illustrated in Fig. 14.
A.4 Comparison of alignment methods
We evaluate the impact of different image alignment strategies on the proposed Doomer dataset, with results summarized in Table 7. Additional global alignment based on homography, implemented via ASpanFormer [8], provides limited improvement, as it fails to handle parallax effects. We also try simply warping the inputs according to the computed optical flow, which is equivalent to applying 1×1 deformable convolution with identity features transform. It substantially improves performance by enabling dense, non-rigid correspondence. However, integrating optical flow into the proposed DCAM consistently yields further gains, demonstrating the advantage of task-aware features learning with respect to alignment. Among all variants, DCAM with PWC-Net achieves the best overall performance, further validating our choice of the optical flow network over the recent of Teed and Deng [33].
Alignment
type
Alignment
network
PSNR, dB ↑ SAM, ° ↓ NSE, % ↓
None None 30.76 4.41 9.62
Homography
AspanFormer [8]
30.44 4.29 9.44
Optical flow-only PWC-Net [32] 30.99 4.10 8.71
RAFT [33] 31.22 4.10 8.24
DCAM PWC-Net [32] 31.46 3.91 8.35
RAFT [33] 31.05 4.53 8.63
Table 7: Impact of different alignment technics on the overall reconstruction quality. Evaluated on the Doomer dataset. Best results are highlighted.